Light travels from air into an optical fiber with an index of refraction of 1.44. (a) In which direction does the light bend? (b) If the angle of incidence on the end of the fiber is 22o, what is the angle of refraction inside the fiber? (c) Sketch the path of light as it changes media.


1
Expert's answer
2021-03-18T13:13:44-0400

Given,

Refractive index of the medium "(\mu)" =1.44

a)

Now, applying the Snell's rule,


"\Rightarrow \frac{\sin i}{\cos r}=\frac{1}{1.44}"


"\Rightarrow \sin i^\circ=\frac{\cos r^\circ}{1.44}"


Hence, ray will move towards the normal normal. Because ray of light moving from rarer medium to denser medium.

b)

Angle of incidence "(i)=22^\circ"

Now, substituting the values in the Snell's equation.

"\sin r=1.44\times \sin 22"

"\Rightarrow r=\sin^{-1}(1.44\times \sin 22^\circ)"

"\Rightarrow r = \sin^{-1}(0.539)"

"\Rightarrow r = 32.61^\circ"

c)